OpenStudy (anonymous):

On a test there are 10 true/false questions. If you guess on each of them, what is the probability that you will get exactly 6 correct?

OpenStudy (anonymous):

It should be \[ \frac {\binom{10} 6}{2^{10}} \]

OpenStudy (anonymous):

It says that I should be using the formula: (nCx)(Px)(1-p)^(n-x) I don't remember what is what.

OpenStudy (anonymous):

Binomial distribution.

jimthompson5910 (jim_thompson5910):

In this case, n = 10, x = 6 and p = 1/2 = 0.5

OpenStudy (anonymous):

How did you get p=1/2?

OpenStudy (campbell_st):

P(true) = 1/2 P(false) = 1/2

jimthompson5910 (jim_thompson5910):

p is the probability of getting an answer choice correct

OpenStudy (anonymous):

Thanks guys